# Improving Targeted Molecule Generation through Language Model Fine-Tuning Via Reinforcement Learning

### v1

@@ Abstract
Developing new drugs is laborious and costly, demanding extensive time investment. In this study, we introduce an innovative de-novo drug design strategy, which harnesses the capabilities of language models to devise targeted drugs for specific proteins. Employing a Reinforcement Learning (RL) framework utilizing Proximal Policy Optimization (PPO), we refine the model to acquire a policy for generating drugs tailored to protein targets. Our method integrates a composite reward function, combining considerations of drug-target interaction and molecular validity. Following RL fine-tuning, our approach demonstrates promising outcomes, yielding notable improvements in molecular validity, interaction efficacy, and critical chemical properties, achieving 65.37 for Quantitative Estimation of Drug-likeness (QED), 321.55 for Molecular Weight (MW), and 4.47 for Octanol-Water Partition Coefficient (logP), respectively. Furthermore, out of the generated drugs, only 0.041% do not exhibit novelty.
Improving Targeted Molecule Generation through Language Model Fine-Tuning Via Reinforcement Learning
Salma J. Ahmed    Mustafa A. Elattar School of Information Technology and Computer Science, Nile University {S.Gamal2121,melattar}@nu.edu.eg
@@ 1 Introduction
The journey from conceptualizing a potential drug to its market availability is lengthy and financially demanding. It must navigate through several critical phases to transform a chemical compound or entity into a viable treatment for human diseases. Initially, a specific molecular target (such as a DNA sequence or protein) associated with a disease must be identified. This target serves as the focal point for drug development, offering the potential for therapeutic intervention. Before a drug can be administered to patients, it must undergo rigorous preclinical research trials. These trials, conducted either in vitro or in vivo, aim to evaluate the drug’s safety profile and assess its effects on biological systems. This phase is pivotal in determining the drug’s potential for therapeutic use and understanding its impact on the body.
Figure 1: MolT5 is initially fine-tuned for compound generation based on input protein. Subsequently, the fine-tuned model is employed in RL-based policy fine-tuning to acquire the ability for targeted compound generation, with drug-target interaction and validity as rewards.
@@
The drug progresses to clinical research after successful preclinical trials, where its efficacy and safety are tested on human subjects. This phase involves carefully designed clinical trials conducted in multiple stages to gather comprehensive data on the drug’s performance and potential side effects. Subsequently, the accumulated data undergoes thorough scrutiny during the FDA review process, where regulatory authorities assess the drug’s safety, efficacy, and overall benefit-risk profile. This extensive journey, while essential for ensuring patient safety and the effectiveness of medications, poses significant challenges. The prolonged timeline and substantial financial investment associated with drug development contribute to the high costs and low success rates in discovering new drugs. As a result, innovative approaches that streamline the drug discovery process and enhance efficiency are crucial for addressing unmet medical needs and advancing therapeutic interventions.
With the rapid advancements in deep learning, many researchers are delving into applying deep learning methodologies to address challenges within the drug discovery domain. These encompass various areas including Drug-Target Interaction Wen et al. (2017); Abbasi et al. (2021); You et al. (2019); Feng et al. (2018); Wang et al. (2020), Protein-Protein Interaction Sun et al. (2017); Hashemifar et al. (2018); Hu et al. (2022); Lee (2023); Peng and Lu (2017), Drug-Drug Interaction Deng et al. (2020); Kumar Shukla et al. (2020); Zhang et al. (2020); Lee et al. (2019); Qiu et al. (2021), De-Novo Drug Design Wang et al. (2022); Bai et al. (2022); Krishnan et al. (2021a); Palazzesi and Pozzan (2022); Krishnan et al. (2021b), and Drug Re-purposing Pan et al. (2022); Issa et al. (2021); Aliper et al. (2016); Lee and Chen (2021); Hooshmand et al. (2021). These approaches leverage the power of deep learning techniques to tackle complex problems in drug discovery, offering promising avenues for accelerating the identification and development of novel therapeutics. This highlights the significant potential of employing deep learning methodologies to expedite the discovery of novel drug candidates, ultimately enhancing our ability to combat emerging diseases. In this study, our primary focus is on harnessing the advancements in deep learning techniques to facilitate drug discovery by proposing an approach for De-Novo drug design. Our main objective centered around developing a generative model capable of taking a molecule target (protein) as input, representing a disease within the human body. The aim was to utilize this model to generate novel molecules (drugs) specifically designed to treat or mitigate the effects of this protein within the body.
@@ 2 Related Work
Popova et al. (2018) introduced ReLeaSE, a technique for producing novel targeted chemical compounds. This method integrates generative and predictive deep neural networks: the generative model is trained to generate new chemical compounds, while the predictive model forecasts desired properties. These models are trained independently and combined using a reinforcement learning approach to guide the generation process. Reinvent a Reinforcement Learning (RL) methodology proposed by Blaschke et al. (2020) for generating novel molecules with specific target interactions by employing RL to guide the generative model. This involves utilizing two Recurrent Neural Networks (RNNs) in an actor-critic setup. The critic, functioning as a prior RNN, preserves previous knowledge of the SMILES. Meanwhile, the actor (Agent) is an identical replica of the Prior or a specialized version that has undergone some preliminary training. The Agent selects actions by sampling a batch of SMILES (S), which are then assessed by the Prior and scored through the scoring function. Gupta et al. (2018) introduced a generative RNN-LSTM model for drug synthesis. Initially, the model undergoes training on a dataset of molecular structures to grasp the syntax and patterns inherent in these representations. Subsequently, the RNN-LSTM model undergoes fine-tuning to skew predictions towards specific molecular targets. This adjustment is achieved by leveraging insights gained from the initial training to tailor the model to particular target molecules. Zhang et al. (2023) conducted training on a generator RNN model using molecular data to familiarize it with the syntax of SMILES for molecular representation, enabling the design of novel and effective small molecules. Subsequently, they developed a drug-target interaction model to serve as a reward in a reinforcement learning framework. This model was employed to steer the generation of the RNN model towards specific properties or targeted molecules. Monteiro et al. (2023) introduced a multi-objective approach for drug synthesis, emphasizing properties and selectivity tailored to biological targets. Their method employs a transformer decoder to create drugs and a transformer encoder to forecast desired properties and refine learning through a feedback loop.
@@ 3 Methods
In this section, we will thoroughly examine the methodology components, spanning from the dataset utilized to the generative model and the elements of the reinforcement learning paradigm.
@@ 3.1 Dataset
In our approach, we utilized BindingDB Liu et al. (2007), a publicly available database encompassing binding affinities of protein-ligand complexes, across all our experiments. The dataset selection specifically targeted proteins serving as drug targets, with their corresponding structural information sourced from the Protein Data Bank Bank (1971). Leveraging this dataset facilitated the achievement of our objectives. Additionally, we implemented a filtering criterion within this dataset to exclude complexes featuring protein amino acid lengths exceeding 500, primarily due to computational resource constraints.
@@ 3.2 Generative Model
An essential aspect of the methodology involves utilizing a generative model, which serves as the vital tool for generating molecular compounds, aligning with the stated objective of the method to generate a molecule drug based on a given protein. We employed MolT5 Edwards et al. (2022), a self-supervised learning framework built upon an encoder-decoder transformer architecture Vaswani et al. (2017). Initially, MolT5 underwent pretraining on a substantial volume of unlabeled molecule compound strings and natural language text. Subsequently, the model underwent fine-tuning on two distinct tasks: molecule captioning and text-based de novo molecule generation. In the first task, the model receives a molecule string prompt and aims to generate a caption, while in the second task, it generates a molecule string based on a provided textual description. Encouraging outcomes were achieved on both tasks, motivating the adoption of the model for our objective. Given a protein amino acid, the model is designed to produce a targeted molecule drug. We utilized the base version of MolT5 (molt5-base) and conducted further fine-tuning employing the protein-ligand complexes sourced from BindingDB to enhance the model’s knowledge for our specific task. The fine-tuning process was evaluated using the BLEU score metric, which measures the similarity between the generated molecule and the actual molecule in the protein-ligand complexes, with the protein serving as input to the model. It is worth noting that our intention was not for something other than for the model to replicate the actual molecule compound within the complex. Instead, it was on instilling in the model the ability to generate. molecule compounds based on a protein string. Following fine-tuning, the model demonstrated proficiency in generating molecules given proteins as inputs. We will integrate the molt5-fine-tuned model into our Reinforcement Learning paradigm to steer the learning process toward generating molecules that target specific proteins.
@@
Figure 2: The left side depicts the methods used for reward calculation, focusing on drug-target interaction (DTI), while the right side illustrates the evaluation of molecule validity.
3.3 Drug Target Interaction (DTI)
As our objective entails enabling the generative model to produce molecule compounds tailored to specific proteins, we sought a robust method to gauge the activity of these molecules toward their protein targets. Consequently, we incorporated a Drug Target Interaction (DTI) model into our methodology, utilizing it as a pivotal component for measuring and characterizing compound-protein binding. For this purpose, we implemented DeepPurpose Huang et al. (2020), a PyTorch Toolkit designed for Molecular Modeling and prediction, particularly in the Drug-Target Interaction (DTI) using deep learning methodologies. This toolkit offers various encodings and representations for molecule compounds and protein amino acid sequences. Subsequently, these representations are inputted into an MLP decoder to generate predictions, manifesting as continuous binding scores or binary outputs. In our experimentation, we explored two distinct encoding methods: utilizing Convolutional Neural Networks (CNNs) on SMILES molecule strings and amino acid sequences and employing transformer encoders on substructure fingerprints. Furthermore, we opted for continuous predictions, explicitly focusing on the median inhibitory concentration (IC50) score. To train the model, we utilized BindingDB-filtered protein-ligand complexes. Upon evaluation, both encoding methods yielded promising results. Consequently, we adopted an ensemble learning approach, averaging the predictions from the two trained models, and selected it as the model for integration into our system.
3.4 Reinforcement Learning Fine-tuning
@@
We implemented the Proximal Policy Optimization (PPO) algorithm Schulman et al. (2017), a widely utilized method for training policies in reinforcement learning tasks Yang and Kasneci (2024); Das et al. (2024); Yau et al. (2024); Klein et al. (2024), within our Reinforcement Learning (RL) framework. Our goal was to leverage this algorithm to guide the model in learning a policy where, upon receiving a protein amino acid sequence as input, the model would generate a molecule compound (drug) tailored to the particular protein. We employed the molt5-base model, which we had previously fine-tuned, as the initial model. We then proceeded to update and assess the model’s performance before and after the fine-tuning process. We employed a combined reward function to optimize the model and align it with the desired performance, which will be elaborated upon in the subsequent subsection.
3.4.1 Rewards
The reward selection holds significant importance within the realm of RL, as it serves as the feedback signal given to an agent, reflecting both its actions and the state of the environment. It serves as a measure of the agent’s performance, indicating its effectiveness in accomplishing tasks. Hence, we opted for a composite reward strategy as illustrated in figure 2 to attain the desired level of performance.
@@ 3.4.1 Rewards
DTI: We integrated the Drug Target Interaction (DTI) model into our RL methodology to serve as the reward mechanism. This decision aligned with our objective of training the model to generate molecules tailored to specific proteins. Consequently, when the model produces a molecule compound unrelated to the protein target, it receives a low reward score. Conversely, it receives a higher score when the generated molecule is targeted to the protein.
@@
Validity: In addition to targeting specific proteins, we aim for the generated molecule to meet chemical validity criteria, ensuring its viability as a potential drug. Hence, we incorporated validity as the second reward within the RL paradigm. To assess validity, we employed the rdkit toolkit Landrum (2013) to determine whether the generated molecule complies with established chemical standards.
@@
4 Experimental Results
This section delves into the comprehensive exploration of our approach, detailing each experimental procedure and its corresponding outcomes. Through systematic refinement of each component, we aimed to optimize performance and achieve superior results.
4.1 Drug Generation
Our initial step involved configuring the generative model to produce molecular compounds when provided with a protein amino acid sequence as input. This is a crucial aspect of our experimental pipeline, as illustrated in the left side of figure 1. Through meticulous fine-tuning of BindingDB complexes, the model demonstrated proficiency in generating molecular compounds given protein inputs. However, our objectives extend beyond mere molecule generation; we aim for targeted compounds tailored to the specific protein. To achieve this, we will proceed with additional fine-tuning iterations to imbue the model with a refined policy for learning protein-targeted molecule generation.
@@ 4.1 Drug Generation
4.2 Reinforcement Learning Fine-tuning Setup
We employed the Transformer Reinforcement Learning (TRL) library von Werra et al. (2020) for fine-tuning our reinforcement learning model, illustrated in 1. This library, built upon the transformers framework, is designed explicitly for refining transformer models through various techniques, including supervised fine-tuning (SFT), Reward Modeling (RM), Direct Preference Optimization (DPO), and Proximal Policy Optimization (PPO). We opted for the PPO method to fine-tune our model due to its recognized efficiency and effectiveness in training complex policies. PPO stands out for its ability to learn with fewer samples than alternative methods. Moreover, its utilization of a trust region optimization approach helps prevent drastic policy changes, ensuring stable learning dynamics. Furthermore, PPO demonstrates scalability, making it suitable for addressing large-scale problems and high-dimensional action spaces. During experimentation, we explored different values for the policy optimization parameters, such as $top_{k}$ and $top_{p}$. Our analysis revealed that the most favorable results were achieved with $top_{k}$ set to 50 and $top_{p}$ to 0.95.

### v2

# Reinforcement Learning for Enhanced Targeted Molecule Generation Via Language Models

@@ Abstract
Developing new drugs is laborious and costly, demanding extensive time investment. In this paper, we introduce a de-novo drug design strategy, which harnesses the capabilities of language models to devise targeted drugs for specific proteins. Employing a Reinforcement Learning (RL) framework utilizing Proximal Policy Optimization (PPO), we refine the model to acquire a policy for generating drugs tailored to protein targets. The proposed method integrates a composite reward function, combining considerations of drug-target interaction and molecular validity. Following RL fine-tuning, the proposed method demonstrates promising outcomes, yielding notable improvements in molecular validity, interaction efficacy, and critical chemical properties, achieving 65.37 for Quantitative Estimation of Drug-likeness (QED), 321.55 for Molecular Weight (MW), and 4.47 for Octanol-Water Partition Coefficient (logP), respectively. Furthermore, out of the generated drugs, only 0.041% do not exhibit novelty.
@@ 1 Introduction
The journey from conceptualizing a potential drug to its market availability is lengthy and financially demanding [22, 6]. It must navigate through several critical phases to transform a chemical compound or entity into a viable treatment for human diseases. Initially, a specific molecular target (such as a DNA sequence or protein) associated with a disease must be identified [13, 34, 2]. This target serves as the focal point for drug development, offering the potential for therapeutic intervention. Before a drug can be administered to patients, it must undergo rigorous preclinical research trials. These trials, conducted either in vitro or in vivo, aim to evaluate the drug’s safety profile and assess its effects on biological systems. This phase is pivotal in determining the drug’s potential for therapeutic use and understanding its impact on the body.
@@
The drug progresses to clinical research after successful pre-clinical trials, where its efficacy and safety are tested on human subjects. This phase involves carefully designed clinical trials conducted in multiple stages to gather comprehensive data on the drug’s performance and potential side effects. Subsequently, the accumulated data undergoes thorough scrutiny during the FDA review process, where regulatory authorities assess the drug’s safety, efficacy, and overall benefit-risk profile. This extensive journey, while essential for ensuring patient safety and the effectiveness of medications, poses significant challenges [45, 17]. The prolonged timeline and substantial financial investment associated with drug development contribute to the high costs and low success rates in discovering new drugs. Manual drug discovery development is a resource-intensive process. For instance, new drugs take an average of 12 and 7 years, respectively, from pre-clinical testing to approval, with development costs for drugs exceeding \$1 billion [14, 16, 48]. As a result, innovative approaches that streamline the drug discovery process and enhance efficiency are crucial for addressing unmet medical needs and advancing therapeutic interventions.
Figure 1: MolT5, a Transformer-based model designed for molecular data, is initially fine-tuned (Stage I) for molecule generation given an input protein. Subsequently, the fine-tuned model undergoes reinforcement learning fine-tuning (Stage II) to enable targeted compound generation, using drug-target interaction and validity as rewards.
With the rapid advancements in Deep Learning (DL), many researchers are exploring its application to address challenges in the drug discovery domain. These efforts span various areas, including Drug-Target Interaction (DTI), where models such as Transformers [49] and convolutional neural networks (CNN) [53] are employed to identify interactions between known drugs and targets. This is a significant challenge in drug repositioning and can significantly accelerate the expensive and time-consuming drug development process [52, 1, 56].
Another key area of focus is Protein-Protein Interactions (PPIs), which are crucial for understanding protein function, disease mechanisms, and therapy design. Many studies have developed deep learning models, such as autoencoders [4], to study sequence-based PPI prediction, achieving promising results [46, 20, 30].
Similarly, Drug-Drug Interaction (DDI) is critical for drug safety surveillance, enabling the effective and safe co-prescription of multiple drugs. Several studies have proposed using DL models to predict interactions in this domain [12, 57, 42].
Another critical area is Drug Re-purposing, where the same drug treats different diseases. To address this, several DL-based approaches, such as graph-based [23] and sequence-based models [9], have been proposed to re-purpose existing drugs for new therapies [37, 19].
De-Novo Drug Design is another key focus, generating novel compounds with desirable pharmacological and physicochemical properties. Various DL approaches, including recurrent neural networks, encoder-decoder architectures, reinforcement learning, and generative adversarial networks, have been applied to this task [51, 27, 28].
These approaches highlight the potential of DL techniques to tackle complex challenges in drug discovery, offering promising avenues for identifying and developing novel therapeutics. This underscores the significant potential of employing deep learning methodologies to expedite the discovery of drug candidates and enhance our ability to combat emerging diseases.
In this paper, our primary focus is leveraging advancements in language models and reinforcement learning techniques to facilitate drug discovery through an approach for De-Novo drug design. Our main objective is to develop a generative model capable of taking a molecular target (protein) as input and using this model to generate novel molecules (drugs) specifically designed to treat or mitigate the effects of this protein within the body. The quality of the generated drugs can be evaluated by measuring their chemical properties to assess solubility and permeability, following [32].
The rest of the paper is organized as follows: Section II provides an overview of the related literature. Section III discusses the components of the proposed methodology. Section IV describes the experimental setup and presents the results. Section V offers an in-depth analysis and discussion of the findings. Finally, Section VI concludes the paper by highlighting key insights and directions for future research.
@@ 2 Related Work
In [40], the authors introduced ReLeaSE, a technique for producing novel targeted chemical compounds. This method integrated generative and predictive deep neural networks, where the generative model was trained to generate new chemical compounds while the predictive model forecasts desired properties. These models are trained independently and combined using a reinforcement learning approach to guide the generation process. Reinvent the Reinforcement Learning (RL) approach described by [8] for generating novel molecules that interact with specific targets by using RL to steer the generative model. This approach employs two Recurrent Neural Networks (RNNs) in an actor-critic setup. The critic, acting as the prior RNN, retains previous knowledge of the SMILES representation. The actor, or agent, is either an identical copy of the prior or a modified version that has undergone some initial training. The agent selects actions by sampling a batch of SMILES (S), which are then evaluated by the prior RNN and scored using a predefined scoring function. The method in [18] introduced a generative RNN-LSTM model for drug synthesis. Initially, the model undergoes training on a dataset of molecular structures to grasp the syntax and patterns inherent in these representations. Subsequently, the RNN-LSTM model undergoes fine-tuning to skew predictions towards specific molecular targets. This adjustment is achieved by leveraging insights gained from the initial training to tailor the model to particular target molecules. Training on a generator RNN model conducted by [58] using molecular data to familiarize it with the syntax of SMILES for molecular representation, enabling the design of novel and effective small molecules. Subsequently, they developed a drug-target interaction model to serve as a reward in a reinforcement learning framework. This model was employed to steer the generation of the RNN model towards specific properties or targeted molecules. A multi-objective approach was introduced by [36] for drug synthesis, emphasizing properties and selectivity tailored to biological targets. Their method employed a transformer decoder to create drugs and a transformer encoder to forecast desired properties and refine learning through a feedback loop.
@@ 3 Methods
In this section, we thoroughly examine the methodology components, spanning from the dataset utilized to the generative model and the elements of the reinforcement learning paradigm.
@@ 3.1 Dataset
In all experiments and components of the proposed method, we use BindingDB [33], a publicly available database containing binding affinities of protein-ligand complexes. The dataset includes target proteins that serve as drug targets, with their corresponding structural information obtained from the Protein Data Bank [5]. It contains about 2,993,668 binding complexes data utilizing 9,499 proteins and over 1,301,732 molecules. This dataset is instrumental in achieving our objectives. Furthermore, a filtering criterion was applied to exclude complexes with protein amino acid lengths exceeding 500, primarily due to computational resource limitations.
@@ 3.2 Generative Model
Figure 2: The lefthand side depicts the methods used for reward calculation, focusing on drug-target interaction, while the righthand side illustrates the evaluation of molecule validity.
An essential aspect of the methodology involves utilizing a generative model, which is vital for generating molecular compounds based on a given protein. We employed MolT5 [15], a self-supervised learning framework built upon an encoder-decoder transformer architecture [49]. MolT5 is pre-trained on unlabeled molecule compound strings and natural language text. Subsequently, the model underwent fine-tuning on two distinct tasks: molecule captioning, where the model receives a molecule string prompt and aims to generate a caption. In contrast, in the second task, text-based de novo molecule generation, the model generates a molecule string based on a provided textual description. Encouraging outcomes were achieved on both tasks, as stated by the authors, motivating the adoption of the model for our objective. Given a protein amino acid, the model should generate a targeted molecule drug. We utilize the base version of MolT5 (molt5-base) and conduct the first stage of the fine-tuning employing the protein-ligand complexes sourced from BindingDB to enhance the model’s knowledge for our specific task, as depicted in Stage I of Figure 1, where the model takes a protein as input and generates a molecule, which is then compared with the true molecule in the protein-ligand complex.
3.3 Drug-Target Interaction
We aim to enable the generative model [15] to generate molecule compounds tailored to specific proteins. To assess the activity of these molecules, we integrated a Drug Target Interaction (DTI) model using DeepPurpose [21], a PyTorch toolkit for molecular modelling and prediction. It provides encodings for molecules and protein sequences fed into a multi-layer perceptron (MLP) decoder to predict binding scores. We tested two encoding methods: CNNs for SMILES strings and protein sequences and a CNN for protein sequences combined with transformer encoders for substructure fingerprints. Focusing on IC50 scores, we trained the models using BindingDB data and adopted an ensemble learning approach, merging predictions from both models.
3.4 Reinforcement Learning Setup
@@
We used the Proximal Policy Optimization (PPO) algorithm [43], a widely utilized method for training policies in reinforcement learning tasks [54, 10, 55, 26], within our Reinforcement Learning (RL) framework. Our goal was to use this algorithm to guide the model in learning a policy where, upon receiving a protein amino acid sequence as input, the model would generate a molecule compound (drug) tailored to the specific protein, as shown in Stage II of Figure 1. We used the molt5-base model, which we had previously fine-tuned, as the starting point. We then fine-tuned it a second time using an RL policy. Reward selection is crucial in RL, as it acts as the feedback signal given to an agent, reflecting its actions and the state of the environment. It measures the agent’s performance and indicates its effectiveness in completing tasks. Therefore, we adopted a composite reward strategy, as illustrated in Figure 2, to achieve the desired level of performance. The rewards used in our approach are as follows:
@@
  • DTI: We integrated the DTI model into our RL methodology to serve as the reward mechanism. This decision aligned with our objective of training the model to generate molecules tailored to specific proteins. Consequently, when the model produces a molecule compound unrelated to the protein target, it receives a low reward score. Conversely, it gets a higher score when the generated molecule is targeted to the protein.
@@
  • Validity: In addition to targeting specific proteins, we aim for the generated molecule to meet chemical validity criteria, ensuring its viability as a potential drug. Hence, we incorporated validity as the second reward within the RL paradigm. To assess validity, we employed the rdkit toolkit [29] to determine whether the generated molecule complies with established chemical standards.
@@
4 Experimental Setup and Results
This section details each experimental procedure and its outcomes. Through systematic refinement of each component, we aimed to optimize performance and achieve superior results.
4.1 Fine-Tuning Stage I
Our initial fine-tuning stage focused on configuring the generative model to produce molecular compounds when provided with a protein amino acid sequence as input. This step represents a critical component of our experimental pipeline, as illustrated in Stage I of Figure 1. The model is fine-tuned on BindingDB complexes and evaluated using the Bilingual Evaluation Understudy (BLEU) metric [38], which measures the similarity between the generated molecule and the actual molecule in the protein-ligand complexes, with the protein serving as input. Following this fine-tuning process, the model demonstrated the ability to create molecules based on protein inputs. However, our objectives extend beyond simple molecule generation; we aim to produce targeted molecules tailored to each protein. We conduct additional RL fine-tuning iterations to enable the model to learn a more refined policy.
4.2 Fine-tuning Stage II
We employed the Transformer Reinforcement Learning (TRL) library [50] for the RL fine-tuning. This library, built upon the transformers framework, is designed explicitly for refining transformer models through various techniques, including supervised fine-tuning (SFT), Reward Modeling (RM), Direct Preference Optimization (DPO), and Proximal Policy Optimization (PPO). We opted for the PPO method to fine-tune our model due to its recognized efficiency and effectiveness in training complex policies. PPO stands out for its ability to learn with fewer samples than alternative methods. Moreover, its utilization of a trust region optimization approach helps prevent drastic policy changes, ensuring stable learning dynamics. During experimentation, we explored different values for the policy optimization parameters, such as $top_{k}$ and $top_{p}$. Our analysis revealed that the most favourable results were achieved with $top_{k}$ set to 50 and $top_{p}$ to 0.95.